# Logarithmic Enriques varieties

### v1

@@ 4. Enriques manifolds and log-Enriques varieties
Inspired by [15, 47, 54], we give the following general definition of Enriques manifolds. We intend to give a sufficiently flexible definition to ensure that there are enough examples to make the theory interesting and to prepare the generalization given below to a singular setup using a similar definition.
@@
Let $X$ be an Enriques manifold and consider its canonical cover $Y\to X$, which is a projective manifold with trivial canonical divisor. By the Beauville-Bogomolov decomposition theorem [7], up to an étale cover, the variety $Y$ splits as a product of IHS manifolds, CY manifolds and an abelian manifold. Since there is in general not much control on this étale cover, we focus on special types of Enriques manifolds where the variety $Y$ itself is split, and we are mostly interested in the “atomic” types:
@@
Our definition is less restrictive than those of [15] since we do not require any condition on the holomorphic Euler characteristic of $X$. It is also more general than that of [47] that are what we call here Enriques manifolds of symplectic type, since every étale quotient of a projective IHS manifold is cyclic. We will prove below that for both these types, it is equivalent to ask that the canonical cover is of the given type. We could a priori define a third abelian type by the condition of the existence of a cyclic étale cover by an abelian manifold, as a special case inside the subfamily of generalized hyperelliptic varieties called Bagnera – de Franchis varieties (see [24]), however we show below that this does not exist.
@@
We use an argument taken from [24, Remark 2.8], communicated to us by Martina Monti. Let $A=V/\Lambda$ be an abelian variety, where $\Lambda$ is a lattice in a finite dimensional complex vector space $V$, and assume that $g\in\mathrm{Aut}(A)$ is an automorphism acting freely such that $A/\langle g\rangle=X$. Decomposing $g(x)=\alpha x+b$ where $\alpha\in\mathrm{GL}(V)$ and $b\in\Lambda$, the freeness is equivalent to the property that there exist no $(x,\ell)\in V\times\Lambda$ such that $(\alpha-\mathrm{id})x=\ell-b$, so $\alpha-\mathrm{id}$ is not invertible. Let $z$ be an eigenvector of $\alpha$ for the eigenvalue $1$. Then $g^{\ast}dz=dz$ and hence $dz$ descends to an element of $H^{0}(X,\Omega^{1}_{X})\cong H^{1}(X,\mathcal{O}_{X})$, so $q(X)>0$: this is a contradiction. ∎
@@
To extend the notion of Enriques manifold to a singular setup, in a similar spirit as above, inspired by [56, Definition 1.1] we give the following definition:
@@
Let $X$ be a log-Enriques variety and consider its canonical cover $Y\to X$. By the decomposition theorem of Höring–Peternell [35], up to a quasi-étale cover, the variety $Y$ splits as a product of ISVs, (singular) CY varieties and an abelian manifold. As in the nonsingular setup, we focus in the ”atomic” types:
@@
In the symplectic setup, if the quasi–étale cover $Y$ is resp. an IHS manifold, an ISV or a PSV, we will make the terminology more precise by saying that $X$ is a log-Enriques variety of IHS type, ISV type or PSV type. Unlike the nonsingular setup, some different types of log-Enriques varieties may well intersect, see §6.2.2.
Remark 4.8.
By [28, Lemma 2.19] and [46, Proposition 2.10], the augmented irregularity of a log-Enriques variety $X$ satisfies the condition $\widetilde{q}(X)\leq\dim X$. If $X$ is of ISV or CY type, then $\widetilde{q}(X)=0$ so log-Enriques varieties of ISV type or of CY type cannot be simultaneously of abelian type.
@@ 5. Log-Enriques varieties of symplectic type
The statement and the argument follow the lines of [47, §2]. Since $p$ is a finite projective morphism, we have $\chi(\mathcal{O}_{Y})=\chi(p_{\ast}\mathcal{O}_{Y})$. Since $p\colon Y\to X$ is an étale cyclic covering of order $d$, there exists a line bundle $L$ on $X$ such that $p_{\ast}\mathcal{O}_{Y}\cong\mathcal{O}_{X}\oplus L\oplus\cdots\oplus L^{d-1}$ with $L^{\otimes d}\cong\mathcal{O}_{X}$. We thus have $\chi(\mathcal{O}_{Y})=\sum_{i=0}^{d-1}\chi(L^{i})$. Since $X$ is nonsingular, by the Hirzebruch–Riemann–Roch theorem, the Euler characteristics of $L^{i}$ depends only on the rational numerical class of $L$, which is trivial since $L^{d}\cong\mathcal{O}_{X}$. We thus get $\chi(\mathcal{O}_{Y})=d\chi(\mathcal{O}_{X})$.
@@
Let $Y$ be an ISV and $\varphi\in\mathrm{Aut}(Y)$ be a symplectic automorphism of order $d\geq 2$ and let $X\coloneqq Y/\langle\varphi\rangle$. Then $\chi(\mathcal{O}_{Y})=\chi(\mathcal{O}_{X})$ and $K_{X}=0$. In particular, a symplectic automorphism of finite order of an IHS manifold is never étale.
@@
The argument is inspired by the proof of [47, Lemma 2.3]. The variety $X$ is normal and projective, with klt  singularities by [40, Proposition 5.20]. We keep notation as in the proof of Proposition 3.2.
Since $\varphi$ is symplectic, the regular parts of the fixed loci of its nontrivial powers are symplectic submanifolds so their codimensions are at least two, hence $\varphi$ is quasi-étale and the quotient $p$ restricts to an étale cyclic covering $p\colon Y^{\circ}\to X_{\rm reg}$. The restriction maps $H^{0}(Y,\Omega^{[i]}_{Y})\to H^{0}(Y^{\circ},\Omega^{i}_{Y^{\circ}})$ are isomorphisms, generated on the right hand side by the restrictions of $\wedge^{i}\sigma_{Y}$ to $Y^{\circ}$ since $Y$ is an ISV. If $\varphi$ is symplectic, all these spaces are $\varphi$-invariant so we get by Hodge symmetry for singular spaces:
@@
In particular, if $Y$ is an IHS and if $\varphi$ is symplectic and étale of order $d\geq 2$, then using Proposition 5.1 we get $\frac{1}{2}\dim Y+1=\chi(\mathcal{O}_{Y})=\chi(\mathcal{O}_{X})=d\chi(\mathcal{O}_{Y})$: this is a contradiction so $\varphi$ cannot be étale. ∎
@@
Let $Y$ be a projective IHS manifold that is deformation equivalent to the Hilbert scheme of $2$ points on a K3 surface. Assume that $Y$ admits a symplectic involution $\iota$. Then the quotient $Y/\iota$ has trivial canonical bundle and it admits a partial resolution that is an ISV (more precisely an irreducible symplectic orbifold), called a Nikulin orbifold [21, Definition 3.1].
@@
Let $Y$ be an ISV of dimension $2n$ with a quasi-étale purely non symplectic automorphism $\varphi$ of order $d$. If $d$ divides $n$ then $K_{X}=0$, otherwise $X\coloneqq Y/\langle\varphi\rangle$ is a log-Enriques variety (of ISV type) and index $r\geq 2$ given as the smallest integer such that $r$ divides $d$ and $d$ divides $rn$.
@@
Let $Y$ be a PSV with symplectic form $\sigma_{Y}$ and let $\varphi$ be a nonsymplectic automorphism of finite order on $Y$. Then any irreducible component of the fixed locus of $\varphi$ in $Y_{\rm reg}$ is isotropic for $\sigma_{Y}$.
@@
Denote the quasi-étale quotient by $p\colon Y\to Y/\langle\varphi\rangle\coloneqq X$. The variety $X$ is normal and projective, with klt  singularities by [40, Proposition 5.20]. By Proposition 3.2(ii), the torsion index $r$ of the Weil divisor $K_{X}$ divides $d$. Using similar arguments as in the proof of Proposition 5.2, we have:
@@
$=H^{0}(Y^{\circ},\omega_{Y^{\circ}})^{\langle\varphi\rangle}=H^{0}(Y^{\circ},\omega_{Y^{\circ}})^{\langle\varphi\rangle}=H^{0}(Y,\omega_{Y})^{\langle\varphi\rangle}.$
The space $H^{0}(Y,\omega_{Y})$ is generated by $\wedge^{n}\sigma_{Y}$. Since $\varphi$ is purely non-symplectic of order $d$, $\varphi^{\ast}\left(\wedge^{n}\sigma_{Y}\right)=\xi^{n}\left(\wedge^{n}\sigma_{Y}\right)$, where $\xi$ is a primitive $d$-th root of unity. If $d$ divides $n$, the space $H^{0}(X,\omega_{X})$ contains a regular global section that does not vanish, so $K_{X}=0$. Otherwise this global section is not $\varphi$-invariant so $H^{0}(X,\omega_{X})=0$ and $K_{X}$ is not trivial. Since $rK_{X}=0$, a similar computation as above gives:
@@
The space $H^{0}(Y,\omega_{Y}^{\otimes r})$ is generated by the nonvanishing section $\left(\wedge^{n}\sigma_{Y}\right)^{\otimes r}$, that is multiplied by $\xi^{rn}$ under the action of $\varphi^{\ast}$, so this global section is $\varphi$-invariant if and only if $d$ divides $rn$. We conclude that $d$ divides $rn$. It follows that $r$ is the smallest integer satisfying both conditions (this argument originates in [47, proof of Proposition 2.8]). ∎
@@
In Proposition 5.4, we see that $r=d$ iff $d$ and $n$ are coprime. In particular, it is not possible to construct an index six (log)-Enriques variety as a quotient of a four-dimensional IHS manifold, see [15, Remark 4.1] and [47, Remark 6.6]. But if $Y$ is an IHS manifold and $\varphi$ is étale, by Proposition 5.1 futhermore $d$ divides $n+1$, so we have always $r=d$.
@@
An Enriques manifold is of symplectic type if and only if its canonical cover is an IHS manifold and it admits a unique cyclic étale IHS cover.
@@
Let $X$ be an Enriques manifold of symplectic type and let $Y\to X$ be a cyclic étale cover of degree $d\geq 2$, where $Y$ is an IHS manifold. Following [47], the cyclic group is generated by a purely nonsymplectic automorphism otherwise some of its powers would be symplectic, and using the holomorphic Lefschetz formula we see that the quotient map would be ramified. By Remark 5.6, the index of the canonical divisor of $X$ is equal to $d$. The canonical covering of $Z\to X$ has thus order $d$ and, since $Y$ is simply connected, $Y$ is isomorphic to $Z$. ∎
@@
Let $S$ be a projective K3 surface with a nonsymplectic involution $\iota$. The natural involution $\iota^{[2]}$ induced on the Hilbert square $S^{[2]}$ of $S$ is nonsymplectic. Its fixed locus is non empty, even if $\iota$ is an Enriques involution. The quotient map $S^{[2]}\to S^{[2]}/\langle\iota^{[2]}\rangle\eqqcolon X$ is quasi-étale and $K_{X}=0$. Similarly as in the proof of Proposition 5.2, we have:
@@
The symplectic form $\sigma_{Y}$ generates $H^{0}(S^{[2]},\Omega_{S^{[2]}}^{[2]})$ and is not invariant, but its exterior product $\wedge^{2}\sigma_{Y}$ is $\varphi$-invariant and generates $H^{0}(S^{[2]},\Omega_{S^{[2]}}^{4})$, so $\chi(\mathcal{O}_{X})=2$ whereas $\chi(\mathcal{O}_{Y})=3$. Camere–Garbagnati–Mongardi [22] prove that $X$ admits a crepant resolution by a Calabi–Yau manifold. By Proposition 3.2, we have an exact sequence:
@@
and $\mathrm{Pic}(S^{[2]})$ is torsion-free. See §6.1.3 for a variant of this contruction producing Enriques manifolds.
@@
Let $X$ be a log-Enriques variety of ISV type of index $d$, and let $p\colon Y\to X$ be a cyclic quasi-étale cover such that $Y$ is an ISV. Assume that $p$ has degree $d$. Then $X$ is the quotient of $Y$ by a purely non symplectic automorphism of order $d$. If moreover $\mathrm{Cl}(Y)$ is torsion-free, then $Y$ is the canonical cover of $X$.
@@
In the above proof, the variety $\overline{Y}$ is not necessarily an ISV, and not even its terminalization. Indeed, if $Y$ is smooth, then by Proposition 5.2 and by [49, Lemma 1], the terminalization of $\overline{Y}$ is an ISV. If $Y$ is a singular ISV, then we would at least need to assume that $Y^{\varphi^{k}}$ is not empty, otherwise $\overline{Y}$ would not be simply connected, but so far no analogue of Proposition 5.2 is known for symplectic actions of finite order on ISVs.
Let us recall some classical basic facts about the topology of a singular quasi-projective complex algebraic variety $X$. When working with the classical topology, $X$ is Hausdorff, connected and locally path-connected and it has the homotopy type of a finite CW-complex, so it is in particular locally simply connected. It thus admits a universal cover (unramified) $p\colon\widetilde{X}\to X$ that inherits a complex analytic structure such that $p$ is a local analytic isomorphism, but in general $\widetilde{X}$ is not algebraic. Assuming that $X$ is normal, it is locally irreducible for the classical topology and it follows that $\widetilde{X}$ is irreducible and that the natural map $\pi_{1}(X_{\rm reg})\to\pi_{1}(X)$ is surjective (see [27, §0.7] and [1, §2] and references therein).
If the cyclic cover $p\colon Y\to X$ defining a log-Enriques variety of ISV type and index $d$ is étale, then clearly $\pi_{1}(X)\cong\mathbb{Z}/d\mathbb{Z}$. Otherwise, the topological fundamental group of $X$ may be smaller than expected. In particular, if the group $\mu_{d}$ acting on $Y$ has a fixed point, then $X$ is simply connected (this is a special case of a result due to Kaneko–Yoshida, see [26, Lemma 1.2]).
Enriques manifolds are defined in [47] as projective, nonsingular varieties, nonsimply connected, whose universal cover is an IHS manifold. In the singular setup, the behavior of the fundamental group and of the universal cover made us choose to use instead the properties of the canonical sheaf to define log-Enriques varieties. However, the following result remains true in the singular setup, but it is not a characterization any more:
@@
Let $X$ be a $K$-torsion variety, non simply connected, whose universal cover $Y$ is an ISV. If the fundamental group $\pi_{1}(X)$ acts purely nonsymplectically on $Y$, then $X$ is a log-Enriques variety of ISV type.
@@
The proof extends [47, Proposition 2.4] to the singular setup, using similar arguments. The fundamental group of $X$ acts biholomorphically on its universal cover $Y$, so $\pi_{1}(X)$ is a subgroup of $\mathrm{Aut}(Y)$. The Lie algebra of $\mathrm{Aut}(Y)$ is isomorphic to $H^{0}(Y,T_{Y})$ (see [42, Lemma 3.4]). Since $Y$ is an ISV, we have $H^{0}(Y,T_{Y})=\{0\}$ by [5, Lemma 4.6], so $\mathrm{Aut}(Y)$ is a discrete group. Since the quotient morphism $p\colon Y\to X=Y/\pi_{1}(X)$ is étale, the group $\pi_{1}(X)$ is in bijection with any fiber of $p$. These fibers are discrete and projective, hence finite, so $\pi_{1}(X)$ is finite. By assumption, the group $\pi_{1}(X)$ acts purely nonsymplectically on $Y$, so the group morphism $\pi_{1}(X)\to\mathbb{C}^{\ast}$ sending an automorphism $\varphi$ to the complex number $\xi$ such that $\varphi^{[\ast]}\omega_{Y}=\xi\omega_{Y}$ is thus injective. We deduce that $\pi_{1}(X)$ is a cyclic group. By Proposition 5.4, since $X$ is assumed to be a $K$-torsion variety, it is a log-Enriques variety. ∎

### v2

@@ 4. Enriques manifolds and log-Enriques varieties
Inspired by [17, 51, 59], we give the following general definition of Enriques manifolds. We intend to give a sufficiently flexible definition to ensure that there are enough examples to make the theory interesting and to prepare the generalization given below to a singular setup using a similar point of view.
@@
Let $X$ be an Enriques manifold and consider its canonical cover $Y\to X$, which is a projective manifold with trivial canonical divisor. By the Beauville-Bogomolov decomposition theorem [7], up to an étale cover, the variety $Y$ splits as a product of IHS manifolds, CY manifolds and an abelian manifold. Since there is in general not much control on this étale cover, we focus on special types of Enriques manifolds where the variety $Y$ itself is split, and we are mostly interested in the “atomic” types.
@@
Our definition is less restrictive than that of [17] since we do not require any condition on the holomorphic Euler characteristic of $X$. It is also more general than that of [51] that is what we call here Enriques manifolds of symplectic type, since every étale quotient of a projective IHS manifold is cyclic. We will prove below that for both types, it is equivalent to ask that the canonical cover is of the given type. We could a priori define a third abelian type by the condition of the existence of a cyclic étale cover by an abelian manifold, as a special case inside the subfamily of generalized hyperelliptic varieties called Bagnera – de Franchis varieties (see [26]), however we show below that this does not exist.
@@
We use an argument taken from [26, Remark 2.8], communicated to us by Martina Monti. Let $A=V/\Lambda$ be an abelian variety, where $\Lambda$ is a lattice in a finite dimensional complex vector space $V$, and assume that $g\in\mathrm{Aut}(A)$ is an automorphism acting freely such that $A/\langle g\rangle=X$. Decomposing $g(x)=\alpha x+b$ where $\alpha\in\mathrm{GL}(V)$ and $b\in\Lambda$, the freeness is equivalent to the property that there exist no $(x,\ell)\in V\times\Lambda$ such that $(\alpha-\mathrm{id})x=\ell-b$, so $\alpha-\mathrm{id}$ is not invertible. Let $z$ be an eigenvector of $\alpha$ for the eigenvalue $1$. Then $g^{\ast}dz=dz$ and hence $dz$ descends to an element of $H^{0}(X,\Omega^{1}_{X})\cong H^{1}(X,\mathcal{O}_{X})$, so $q(X)>0$: this is a contradiction. ∎
@@
To extend the notion of Enriques manifold to a singular setup, in a similar spirit as above, inspired by [61, Definition 1.1] we give the following definition:
@@
Let $X$ be a log-Enriques variety and consider its canonical cover $Y\to X$. By the decomposition theorem of Höring–Peternell [39], up to a quasi-étale cover, the variety $Y$ splits as a product of ISVs, (singular) CY varieties and an abelian manifold. As in the nonsingular setup, we focus on the ”atomic” types:
@@
In the symplectic setup, if the quasi–étale cover $Y$ is respectively an IHS manifold, an ISV or a PSV, we will make the terminology more precise by saying that $X$ is a log-Enriques variety of respectively IHS type, ISV type or PSV type.
Lemma 4.8.
Log-Enriques varieties of ISV, CY or abelian types exclude each other.
Proof.
The argument has been communicated to us by Nikolaos Tsakanikas. By [32, Lemma 2.19] and [50, Proposition 2.10], the augmented irregularity of a log-Enriques variety $X$ is finite (precisely $\widetilde{q}(X)\leq\dim X$) and it is invariant under quasi-étale covers. If $X$ is of ISV or CY type, then $\widetilde{q}(X)=0$ so log-Enriques varieties of ISV type or of CY type cannot be simultaneously of abelian type. Moreover, by [32, Proposition F] the restricted holonomy of a CY variety is a special unitary group whereas those of an ISV is a symplectic group, and by [32, Lemma 4.7] the restricted holonomy is invariant under quasi-étale covers, so the ISV types and the CY types exclude each other. ∎
This result is similar to the nonsingular setup, but log-Enriques varieties of PSV type may well be of some other type too. In a different spirit, we give in §6.2.2 an example of a log-Enriques variety of IHS types that is a quasi-étale cover of a weak Calabi–Yau variety.
@@ 5. Log-Enriques varieties of symplectic type
The statement and the argument follow [51, §2]. Since $p$ is a finite projective morphism, we have $\chi(\mathcal{O}_{Y})=\chi(p_{\ast}\mathcal{O}_{Y})$. Since $p\colon Y\to X$ is an étale cyclic cover of order $d$, there exists a line bundle $L$ on $X$ such that $p_{\ast}\mathcal{O}_{Y}\cong\mathcal{O}_{X}\oplus L\oplus\cdots\oplus L^{d-1}$ with $L^{\otimes d}\cong\mathcal{O}_{X}$. We thus have $\chi(\mathcal{O}_{Y})=\sum_{i=0}^{d-1}\chi(L^{i})$. Since $X$ is nonsingular, by the Hirzebruch–Riemann–Roch theorem, the Euler characteristic of $L^{i}$ depends only on the rational numerical class of $L$, which is trivial since $L^{d}\cong\mathcal{O}_{X}$. We thus get $\chi(\mathcal{O}_{Y})=d\chi(\mathcal{O}_{X})$.
@@
Let $Y$ be a PSV and $\varphi\in\mathrm{Aut}(Y)$ be a symplectic automorphism of order $d\geq 2$ and let $X\coloneqq Y/\langle\varphi\rangle$. Then $\chi(\mathcal{O}_{Y})=\chi(\mathcal{O}_{X})$ and $K_{X}=0$. In particular, a symplectic automorphism of finite order of an IHS manifold is never étale.
@@
The argument is inspired by the proof of [51, Lemma 2.3]. The variety $X$ is normal and projective, with klt  singularities by [44, Proposition 5.20]. We keep notation as in the proof of Proposition 3.2.
Since $\varphi$ is symplectic, the regular parts of the fixed loci of its nontrivial powers are symplectic submanifolds so their codimensions are at least two, hence $\varphi$ is quasi-étale and the quotient $p$ restricts to an étale cyclic cover $p\colon Y^{\circ}\to X_{\rm reg}$. The restriction maps $H^{0}(Y,\Omega^{[i]}_{Y})\to H^{0}(Y^{\circ},\Omega^{i}_{Y^{\circ}})$ are isomorphisms, generated on the right hand side by the restrictions of $\wedge^{i}\sigma_{Y}$ to $Y^{\circ}$ since $Y$ is a PSV. If $\varphi$ is symplectic, all these spaces are $\varphi$-invariant so we get by Hodge symmetry for singular spaces:
@@
In particular, if $Y$ is an IHS manifold and if $\varphi$ is symplectic and étale of order $d\geq 2$, then using Proposition 5.1 we get $\frac{1}{2}\dim Y+1=\chi(\mathcal{O}_{Y})=\chi(\mathcal{O}_{X})=d\chi(\mathcal{O}_{Y})$: this is a contradiction so $\varphi$ cannot be étale. ∎
@@
Let $Y$ be a projective IHS manifold that is deformation equivalent to the Hilbert scheme of two points on a K3 surface. Assume that $Y$ admits a symplectic involution $\iota$. Then the quotient $Y/\iota$ has trivial canonical bundle and it admits a partial resolution that is an ISV (more precisely an irreducible symplectic orbifold), called a Nikulin orbifold [23, Definition 3.1].
@@
Let $Y$ be a PSV (resp. ISV) of dimension $2n$ with a quasi-étale purely non symplectic automorphism $\varphi$ of order $d$. If $d$ divides $n$ then $K_{X}=0$, otherwise $X\coloneqq Y/\langle\varphi\rangle$ is a log-Enriques variety of PSV type (resp. ISV type), and index $r\geq 2$ given as the smallest integer such that $r$ divides $d$ and $d$ divides $rn$.
@@
Let $Y$ be a PSV of dimension at least $4$ with symplectic form $\sigma_{Y}$ and let $\varphi$ be a nonsymplectic automorphism of finite order on $Y$. Then any irreducible component of the fixed locus of $\varphi$ in $Y_{\rm reg}$ is isotropic for $\sigma_{Y}$.
@@
Denote the quasi-étale quotient by $p\colon Y\to Y/\langle\varphi\rangle\coloneqq X$. The variety $X$ is normal and projective, with klt  singularities by [44, Proposition 5.20]. By Proposition 3.2(ii), the torsion index $r$ of the Weil divisor $K_{X}$ divides $d$. Using similar arguments as in the proof of Proposition 5.2, we have:
@@
$=H^{0}(Y^{\circ},\omega_{Y^{\circ}})^{\langle\varphi\rangle}=H^{0}(Y,\omega_{Y})^{\langle\varphi\rangle}.$
The space $H^{0}(Y,\omega_{Y})$ is generated by $\wedge^{n}\sigma_{Y}$. Since $\varphi$ is purely nonsymplectic of order $d$, we have $\varphi^{\ast}\left(\wedge^{n}\sigma_{Y}\right)=\xi^{n}\left(\wedge^{n}\sigma_{Y}\right)$, where $\xi$ is a primitive $d$-th root of unity. If $d$ divides $n$, the space $H^{0}(X,\omega_{X})$ contains a regular global section that does not vanish, so $K_{X}=0$. Otherwise this global section is not $\varphi$-invariant so ${H^{0}(X,\omega_{X})=0}$ and $K_{X}$ is not trivial. Since $rK_{X}=0$, a similar computation as above gives:
@@
The space $H^{0}(Y,\omega_{Y}^{\otimes r})$ is generated by the nonvanishing section $\left(\wedge^{n}\sigma_{Y}\right)^{\otimes r}$, that is multiplied by $\xi^{rn}$ under the action of $\varphi^{\ast}$, so this global section is $\varphi$-invariant if and only if $d$ divides $rn$. We conclude that $d$ divides $rn$. It follows that $r$ is the smallest integer satisfying both conditions (this argument originates in [51, proof of Proposition 2.8]). ∎
@@
We add two observations to Proposition 5.4:
  1. (1)
    We have $r=d$ if and only if $d$ and $n$ are coprime. In particular, it is not possible to construct an index six (log)-Enriques variety as a quotient of a four-dimensional IHS manifold by an order $6$ automorphism, see [17, Remark 4.1] and [51, Remark 6.6], but see §6.2.3 and §6.2.4 for examples of log-Enriques varieties of index $6$ or $12$ using higher-dimensional quotients.
  2. (2)
    If $Y$ is an IHS manifold and $\varphi$ is étale, by Proposition 5.1 futhermore $d$ divides $n+1$, so we have always $r=d$.
@@
An Enriques manifold is of symplectic type if and only if its canonical cover is an IHS manifold, and it admits a unique cyclic étale IHS cover.
@@
Let $X$ be an Enriques manifold of symplectic type and let $Y\to X$ be a cyclic étale cover of degree $d\geq 2$, where $Y$ is an IHS manifold. Following [51], the cyclic group is generated by a purely nonsymplectic automorphism, otherwise some of its powers would be symplectic, and using the holomorphic Lefschetz formula we see that the quotient map would be ramified. By Remark 5.6, the index of the canonical divisor of $X$ is equal to $d$. The canonical cover of $Z\to X$ has thus order $d$ and, since $Y$ is simply connected, $Y$ is isomorphic to $Z$. ∎
@@
Let $S$ be a projective K3 surface with a nonsymplectic involution $\iota$. The natural involution $\iota^{[2]}$ induced on the Hilbert square $Y:=S^{[2]}$ of $S$ is nonsymplectic. Its fixed locus is non empty, even if $\iota$ is an Enriques involution. The quotient map $S^{[2]}\to S^{[2]}/\langle\iota^{[2]}\rangle\eqqcolon X$ is quasi-étale and $K_{X}=0$. Similarly as in the proof of Proposition 5.2, we have:
@@
The symplectic form $\sigma_{Y}$ generates $H^{0}(S^{[2]},\Omega_{S^{[2]}}^{[2]})$ and is not invariant, but its exterior product $\wedge^{2}\sigma_{Y}$ is $\varphi$-invariant and generates $H^{0}(S^{[2]},\Omega_{S^{[2]}}^{4})$, so $\chi(\mathcal{O}_{X})=2$ whereas $\chi(\mathcal{O}_{Y})=3$. Camere–Garbagnati–Mongardi [24] prove that $X$ admits a crepant resolution by a weak Calabi–Yau manifold. By Proposition 3.2, we have an exact sequence:
@@
and $\mathrm{Pic}(S^{[2]})$ is torsion-free. See §6.1.3 for a variant of this construction producing Enriques manifolds.
@@
Let $X$ be a log-Enriques variety of PSV (resp. ISV) type of index $d$, and let $p\colon Y\to X$ be a cyclic quasi-étale cover such that $Y$ is a PSV (resp. ISV). Assume that $p$ has degree $d$. Then $X$ is the quotient of $Y$ by a purely non symplectic automorphism of order $d$. If moreover $\mathrm{Cl}(Y)$ is torsion-free, then $Y$ is the canonical cover of $X$.
@@
By a recent result of Bertini, Grossi, Mauri and Mazzon [11, Proposition 3.16] quotients of ISV by finite order automorphisms acting symplectically are again ISV, so in our proof, if $Y$ is an ISV then $\overline{Y}$ is an ISV. In particular since $Y$ and $\overline{Y}$ are simply connected, the quotient by $\varphi^{k}$ is quasi-étale, but certainly not étale. This remark holds for every quotient of an ISV by a finite order automorphism acting symplectically, so that the second assertion of Proposition 5.2 extends to the singular setting: a symplectic automorphism of finite order of an ISV is never étale.
Let us recall some classical basic facts about the topology of a singular quasi-projective complex algebraic variety $X$. When working with the classical topology, $X$ is Hausdorff, connected and locally path-connected and it has the homotopy type of a finite CW-complex, so it is in particular locally simply connected. Thus it admits a universal cover (unramified) $p\colon\widetilde{X}\to X$ that inherits a complex analytic structure such that $p$ is a local analytic isomorphism, but in general $\widetilde{X}$ is not algebraic. Assuming that $X$ is normal, it is locally irreducible for the classical topology: it follows that $\widetilde{X}$ is irreducible and that the natural map $\pi_{1}(X_{\rm reg})\to\pi_{1}(X)$ is surjective (see [31, §0.7] and [1, §2] and references therein).
If the cyclic cover $p\colon Y\to X$ defining a log-Enriques variety of ISV type and index $d$ is étale, then clearly $\pi_{1}(X)\cong\mathbb{Z}/d\mathbb{Z}$. Otherwise, the topological fundamental group of $X$ may be smaller than expected. In particular, if the group $\mu_{d}$ acting on $Y$ has a fixed point, then $X$ is simply connected (this is a special case of a result due to Kaneko–Yoshida, see [30, Lemma 1.2]).
Enriques manifolds are defined in [51] as projective, nonsingular varieties, nonsimply connected, whose universal cover is an IHS manifold. In the singular setup, the behavior of the fundamental group and of the universal cover pushed us to use instead the properties of the canonical sheaf to define log-Enriques varieties. However, the following result remains true in the singular setup, but it is not a characterization any more.
@@
Let $X$ be a $K$-torsion variety, non simply connected, whose universal cover $Y$ is a PSV (respectively and ISV). If the fundamental group $\pi_{1}(X)$ acts purely nonsymplectically on $Y$, then $X$ is a log-Enriques variety of PSV (respectively ISV) type.
@@
The proof extends [51, Proposition 2.4] to the singular setup, using similar arguments. The fundamental group of $X$ acts biholomorphically on its universal cover $Y$, so $\pi_{1}(X)$ is a subgroup of $\mathrm{Aut}(Y)$. The Lie algebra of $\mathrm{Aut}(Y)$ is isomorphic to $H^{0}(Y,T_{Y})$ (see [46, Lemma 3.4]). Since $Y$ is an ISV, we have $H^{0}(Y,T_{Y})=\{0\}$ by [5, Lemma 4.6], so $\mathrm{Aut}(Y)$ is a discrete group. Since the quotient morphism $p\colon Y\to X=Y/\pi_{1}(X)$ is étale, the group $\pi_{1}(X)$ is in bijection with any fiber of $p$. These fibers are discrete and projective, hence finite, so $\pi_{1}(X)$ is finite. By assumption, the group $\pi_{1}(X)$ acts purely nonsymplectically on $Y$, so the group morphism $\pi_{1}(X)\to\mathbb{C}^{\ast}$ sending an automorphism $\varphi$ to the complex number $\xi$ such that $\varphi^{[\ast]}\omega_{Y}=\xi\omega_{Y}$ is injective. We deduce that $\pi_{1}(X)$ is a cyclic group. By Proposition 5.4, since $X$ is assumed to be a $K$-torsion variety, it is a log-Enriques variety. ∎